Reduce the following equations to standard form, and then identify which conicoids they represent. Further, give a rough sketch of each.
i) x^2 +y^2 +2x−y−z+3 = 0
ii) 3y^2 +3z^2 +4x+3y+z = 9
1
Expert's answer
2020-03-19T17:02:14-0400

i) "x^2+y^2+2x-y-z+3=0"

then we group variables. We get equation. "(x+1)^2+(y-1\/2)^2+7\/4=z"

Finish formul is "(x+1)^2+(y-1\/2)^2=z-7\/4"

This formula is the formula of elliptic paraboloid (we then can see from standard form of elliptic paraboloid, "z-z_0=A(x-x_0)^2+B(y-y_0)^2" , where the coefficients A and B have the same sings )

Picture



ii) "3y^2 +3z^2 +4x+3y+z = 9"

Then, also we do group variables. We get the formula. "3(y+1\/2)^2-3\/4+3(z+1\/6)^2-1\/12=9-4x""3(y+1\/2)^2+3(z+1\/6)^2=49\/6-4x" All numbers and variable x on the right side.

This is also elliptic paraboloid, but this elliptic paraboloid going on x-os.

Picture